let the function p(x) = x6 + ax55 + bx4 +cx3 +dx2 +ex +f be a real polynomial. given leading coefficient is 1 and p(1) = 2,  p(2) = 5, p(3) = 10, p(4) = 17, p(5) = 26 and p(6) = 37 then p(10) = ?

Grade:12th pass

1 Answers

Nishant Vora IIT Patna
7 years ago
Let Q(x) = P(x) – (x^2 + 1)

so clearly x=1,2,3,4,5 and 6 are roots of Q(x)

P(x) – (x^2 + 1) = (x-1)(x-2)(x-3)(x-4)(x-5)(x-6)

hence P(10) = 9*8*7*6*5*4 + (10^2 + 1) = 60581

Thanks